# Kernel Three Pass Regression Filter

### v1

@@ 3. Run time series regression of $y_{t+h}$ on predictive factors $\hat{\boldsymbol{F}}_{t}$,
    $\mathbb{E}\left|M^{-1/2}T^{-1/2}\sum_{s=1}^{T}\sum_{i=1}^{N}\left[\varepsilon_{is}\varepsilon_{it}-\sigma_{ii,st}\right]\right|^{4}\leq R$
@@
5 outlines the utility of using proxies. Proxies are target-relevant in the sense that they only load on the factor(s) which have any explanatory power for the target. Non-singularity of $\boldsymbol{\Lambda}_{f}$ ensures that none of the proxies are redundant.
@@ 5 Empirical Applications
In this section, we apply our proposed method to real-world applications, focusing on forecasting time series variables across various economic domains such as macroeconomics, finance, labor, housing, and prices. To assess the performance of our approach, we conduct comparative analyses against competitive methods, employing the out-of-sample $R^{2}$ performance metric as a benchmark. Additionally, we provide detailed explanations of our data, including the transformations applied to ensure stationarity and the optimization of hyperparameters related to kernel functions.
We compare our method against six different forecasting methods. The first is the PCA-based factor model proposed by Stock & Watson, (2002b); we write it as PCA in our performance tables. Following on PCA-based methods, we compare against PC-Squared (PC-Sq) and Squared-PC(Sq-PC) of Bai & Ng, (2008). The fourth method is the kernel-based PCA (kPCA) method implemented in Kutateladze, (2022). The fifth is our linear counterpart, the 3PRF. The last is an autoregressive model of lag order two. Some of these methods require tuning of hyper-parameters to provide the best results, we do tune them as discussed in the subsection-5.2. Further, kernel method-based papers report results with different kernels; however, we only report the results with a Gaussian kernel to save some space.
5.1 Performance Metric: Out of Sample $R^{2}$
@@
It is noteworthy that out-of-sample $R^{2}$ ranges from $\infty$- ∞ to $1$, unlike in-sample $R^{2}$, which ranges from zero to one. A positive out-of-sample $R^{2}$ indicates that the forecasting method outperforms the historical average. At the same time, a negative value suggests that the forecasting method performs worse than a simple method that forecasts $y_{i}$ equal to the historical average.
We adopt the rolling window method to compute out-of-sample $R^{2}$, consistent with standard practices in the literature. Appendix B.1 provides a detailed exposition of our methodology.
@@ 5.2 Data and Hyper-Parameter Tuning
We utilize the quarterly macroeconomic dataset, FRED-QD. It covers the time span of 1959-2023. This dataset encompasses a comprehensive array of variables exceeding 250, including macroeconomic (such as GDP, Consumption, and Investment), financial, labor market, housing, and industrial and manufacturing variables. Our forecasting endeavors focus on a single target time series while leveraging the remaining variables within the predictor set. Our main text shows the forecasting of important variables from various domains, such as macro, labor, housing, finance, prices, etc. For ease of reference, we present a tabulation comprising the mnemonic codes and details of the variables in the FRED-QD dataset alongside their counterparts in the Stock-Watson dataset in Appendix B.2.
@@ 5.2.1 Data Transformation
PCA encounters challenges when applied to nonstationary time series commonly found in economics and finance. Nonstationary variables lack a defined population mean, and the sample standard deviation tends to diverge as the number of observations increases. Onatski & Wang, (2021) discusses some of the significant issues arising from applying PCA to nonstationary data in detail. Generally, researchers address this by manually examining each series to identify necessary transformations before computing principal components. Hamilton & Xi, (2024) offers an improved method of transformation to achieve the stationarity of the predictor set. We use their method to remove nonstationarity from our data.
Scholars in the literature often employ sample periods devoid of structural breaks; for instance, Fan et al., (2023) says “There exist significant structural breaks for many variables around the year of financial crisis in 2008 which makes our data non-stationary even after performing the suggested transformations”. Therefore, our study focuses on the stationary period spanning from 1965 to 2007888Another indirect advantage of the choice of this sample period is that it gives us the number of samples less than the number of predictors ($T<N$), hence a truly high-dimensional scenario to test our method in.. For the sanity check, we conduct analyses on different combinations of the sample periods, including the entire available sample period from 1959 to 2023, and find no qualitative discrepancies in our findings.
In our main analysis, the sample period from 1964:Q4 to 2007:Q1 comprises $T=170$ observations (periods) and $N=176$ variables (predictors). While the data is initially available for around 250 series, those with missing values are removed; this leaves us with a total of 176 series. We adopt the rolling window methodology for model training and hyper-parameter tuning, utilizing 70% of the total observations as the width of the rolling window. We observe qualitatively similar performance across varying window sizes (50%, 60% of total data). Subsequently, we evaluate the forecasting performance on out-of-sample data and present the results in the tables.
@@ 5.2.2 Hyper-parameter Tuning
Our methodology incorporates the kernel as a fundamental element of the estimation process. However, the kernel function includes a hyperparameter that necessitates optimization. Concurrently, a similar hyperparameter requires tuning in the context of a competitor method, namely kernel PCA. Thus, we employ an identical tuning procedure for both methodologies. We adopt a Gaussian kernel, which relies on a single hyperparameter, denoted as $\sigma$, for our specific applications. Due to the nature of our dataset, which comprises time series data, conventional methods such as random sampling for k-fold cross-validation are not directly applicable. Instead, we partition the data into two folds and conduct cross-validation to determine the optimal tuning parameter. Further elaboration on the algorithm employed for this purpose is provided in the appendix-B.3.
Furthermore, among our competitive methodologies, certain approaches are based on PCA, necessitating the judicious selection of the number of principal components to be included in the analysis. To address this, we employ the eigenvalue ratio test method proposed by Ahn & Horenstein, (2013). This method computes the ratio of each eigenvalue to its predecessor and selects the number of principal components corresponding to the index where this ratio attains its maximum value.
@@ 5.3 Forecasting Using Theory Guided Proxies
We show the utility of theory-guided proxies as in Kelly & Pruitt, (2015) to predict certain variables. In this section, we present a comparison with the linear benchmark, i.e., 3PRF.
The primary strength of our methodology lies in its ability to accurately forecast over longer horizons, where non-linear relationships become more pronounced. However, for the purposes of this section, we restrict our analysis to one-period ahead forecasting. Our results show that our approach outperforms its closest competitor even in the short term. It is essential to note that the primary objective of this subsection is to establish the viability of theory-guided proxies rather than to conduct a comprehensive comparison with competing methods. A more extensive performance evaluation will be presented in subsequent subsections, where we used the auto-proxy method discussed in table-3 to construct forecasts using Kernel 3PRF.
5.3.1 National Income Identity
Sometimes, the context of the problem gives us good choices for proxy variables. There is ample evidence that consumption and investments can be reasonable proxies for a country’s gross domestic product (GDP). The two proxies affect the GDP through the multiplier effect. Therefore, for the GDP (Y), we can choose consumption (C) and Investment (I) as proxy variables. We forecast GDP by taking combinations of consumption and investment as proxies. The results are presented in Table-4.
Table 4: One-period Ahead Out-of-Sample $R^{2}$ for National Income
@@ 5.3.1 National Income Identity
It is clear that using the proxies works, and our method performs better than the nearest competitive method Kelly & Pruitt, (2015).
5.3.2 Quantity Theory of Money
In another application of theory-guided proxies, we utilize the quantity theory of money to forecast inflation using GDP and money supply as proxies. The quantity theory of money equation states:
@@ 5.3.2 Quantity Theory of Money
Our objective is to forecast inflation. In our dataset, Δ(Real $\Delta(\text{Real Product})$ represents the change in GDP, while Δ(Money $\Delta(\text{Money Supply})$ signifies the changes in money supply using log growth in M1. Similar to Fama (1981), changes in velocity, which are challenging to quantify, serve as the error term in the forecasting relationship. The timing is aligned such that proxies observed at time $t$ are utilized to extract information from the predictors at time $t$ for forecasting inflation at time $t+1$. The results for one-period ahead inflation forecasts are presented in Table-5.
Table 5: One-period Ahead Out-of-Sample $R^{2}$ for Inflation
@@
The results indicate that the theory-guided proxies effectively capture inflation dynamics, yielding performance comparable to the closest competitor. It is important to emphasize again that this analysis focuses on one-step-ahead forecasts, which are not the primary strength of our methodology. The purpose of presenting these results is solely to demonstrate the efficacy of the proxies.
5.4 Long and Short-Run Forecasting: Comparative Performance Plots
Linear equations may approximate the relationships between targets and predictors effectively when the analysis time span is sufficiently short. However, as the time span expands, incorporating non-linearities becomes crucial, as noted by Goulet Coulombe et al., (2022). Linear methods exhibit deteriorating performance as the forecasting horizon extends into the future. Thus, the primary strength of our methodology lies in addressing longer-horizon forecasting challenges. Nevertheless, we demonstrate that even in shorter-horizon forecasting scenarios, our approach distinctly outperforms the nearest competitor Kelly & Pruitt, (2015). In the subsequent section, we present comparative plots, reserving detailed and comprehensive comparisons for subsequent sub-sections.
5.4.1 Short Horizon Forecasting
While kernel functions are recognized for their ability to effectively capture non-linear relationships between variables, it’s essential to acknowledge that linear relationships can also be effectively modeled by kernel functions by selecting an appropriate tuning hyperparameter.
In this sub-section, we aim to demonstrate the superiority of our kernel-based forecasting method over the closest linear forecasting method (3PRF) in the one-step-ahead forecasting problem. We present the one-step-ahead forecasted series using our method (k3PRF) and 3PRF alongside the true values of the target variables in figure-3. Specifically, we compare the performance across four different types of economic series from different domains: macro series Exports, price series GDP Deflator, manufacturing series Industrial Production, and financial series the S&P 500 Index.
@@ 5.4.1 Short Horizon Forecasting
The plots in figure-3 clearly illustrate that our predictions closely align with the true series, surpassing the predictions generated by 3PRF. Our method captures significant fluctuations, including notable drops and peaks, demonstrating its efficacy in short-term forecasting scenarios.
5.4.2 Long Horizon Forecasting
In this section, we plot the long-horizon (twelve periods ahead) forecasts generated by the 3PRF method and our method against the actual value of the target series in figure-4. We compare the two methods on the same four economic series as in the previous subsection.
@@ 5.4.2 Long Horizon Forecasting
We observe that the long-horizon forecasts of our competitor, the 3PRF method, tend to flatter, i.e., approaching a historical mean. Whereas our method can accurately track the true series by capturing ups and downs. Although it is acknowledged that absolute predictability may diminish over longer horizons relative to linear methods, our approach continues to demonstrate robust performance, even on extended forecasting horizons.
5.5 Forecasting Macroeconomic Variables
@@
Results and Disucssion
To present the results in a tidy manner, we make two tables. In table-6, we present the forecasting performance on three series: GDP, Consumption, and Investment. To keep track of the presentation, we informally call it ‘Group-I’. Table-7 shows comparative forecasting performance on ‘Group-II’ macro variables: Exports, Imports, Fixed Investments, and Industrial Production (Final Index). Further detail on these variable series is available in appendix-B.2. The numbers reported in the table are out of sample $R^{2}$ defined earlier in the text. We present performance for various forecast horizons ranging from one period ahead to twelve periods ahead.
An empirical observation highlights that among various unsupervised forecasting methodologies, including PCA, Squared-PC, PC-squared, and non-linear unsupervised approaches such as kernel PCA, none exhibit superior performance compared to our proposed method across any forecast horizon for the seven series under consideration. While the supervised linear forecasting model 3PRF demonstrates improved performance relative to the unsupervised techniques, it still falls short of outperforming our non-linear supervised approach. Notably, the autoregressive (AR) model emerges as the sole contender capable of surpassing our method in the short term. Remarkably, our approach exhibits only a marginal performance differential compared to the AR model in shorter forecast horizons yet significantly outperforms it over longer forecast horizons.
In summary, the consistent empirical observation across all series suggests that our method progressively outperforms all alternative methodologies with increasing forecast horizons. Notably, the AR model remains a competitive rival in the short term, attributed to the linear model’s relative advantage in capturing short-term dynamics where nonlinearities may not yet manifest. Consequently, our method emerges as a dependable and preferred forecasting framework across all forecast horizons in macroeconomic prediction tasks.
5.6 Forecasting Labor Market Variables
The examination of the unemployment rate and total non-farm employment offers a deeper understanding of the labor market dynamics within an economy. The labor market is a pivotal element given its centrality in providing income and livelihoods for most economic actors. Consequently, fluctuations in the labor market exert profound impacts on consumer spending patterns and inflationary pressures. These indicators serve as crucial inputs into important economic decisions, notably informing the formulation of both fiscal and monetary policies to achieve goals. We forecast these two series and present the results in table-8.
Employment
@@ 5.6 Forecasting Labor Market Variables
Table 8: Out of Sample $R^{2}$ of Employment (non-farm) and Unemployment Rate
Result and Discussions
We present the results in table-8. It’s worth noting that the overall trends in the total non-farm employment series align with our macroeconomic time series discussion. Importantly, our method consistently outperforms competitors in longer horizon forecasting ($h\geq 4$), a testament to its robustness and reliability. As we extend the forecast horizons, the margin of our method’s superior performance continues to grow, instilling confidence in its effectiveness. In the short term, our method may not outperform the autoregressive model, but it still ranks second and is only slightly behind. Similarly, in forecasting the unemployment rate, our method consistently outperforms all other methods, including the autoregressive (AR) model, from $h=2$ onwards.
To summarize, our findings mirror those of the macroeconomic series forecasting. Our method consistently demonstrates superior performance across various forecast horizons and labor market indicators, reaffirming its reliability and preference in macroeconomic prediction tasks.
5.7 Forecasting Housing Variables
The housing market is characterized by its multidimensional aspect, involving various natures and functions. Houses serve as places of residence and vehicles for wealth accumulation, asset transfer to the next generation, and investment. Also, housing comprises a significant portion of household assets, making it an important factor in household decision-making. Consequently, distortions in the housing market can affect household spending and the broader economy through institutions such as mortgage banks. Therefore, forecasting housing market activities holds substantial value. In this section, we forecast two housing series: the first is Privately Owned Housing Starts (HStart), and the second is Privately Owned Housing Starts in the Western Census region(HStart-W).
Table 9: Out of Sample $R^{2}$ of Privately Owned Housing Starts: Total and West Region

### v3

@@ 3. Run time series regression of $y_{t+h}$ on predictive factors $\hat{\boldsymbol{F}}_{t}$,
    $\mathbb{E}\left|M^{-1/2}T^{-1/2}\sum_{s=1}^{T}\sum_{i=1}^{M}\left[\varepsilon_{is}\varepsilon_{it}-\sigma_{ii,st}\right]\right|^{4}\leq R$
@@
Assumption 5 outlines the utility of using proxies. Proxies are target-relevant in the sense that they only load on the factor(s) that have any explanatory power for the target. Non-singularity of $\boldsymbol{\Lambda}_{f}$ ensures that none of the proxies are redundant.
@@ 4 Results
We show that our estimated forecast converges to the infeasible best in probability. To show the same, we prove some intermediate results. All the proofs are in the appendix.
@@ 5 Empirical Applications
We apply our proposed method to real-world applications, focusing on forecasting time series variables across various economic domains such as national income, finance, labor, housing, prices, etc. To assess the performance of our approach, we conduct comparative analyses against competitive methods, employing the out-of-sample $R^{2}$ performance metric as a benchmark. Out of sample $R^{2}$ is computed as:
@@
It computes the out-of-sample proximity of our forecast $\hat{y}$ with the target ($y$) relative to a historical mean $(\bar{y})$; a positive value indicates that the forecast is better than the historical mean. Detailed explanations of performance metrics computation are provided in the Supplementary appendix-B.2.
We compare our method against six different forecasting methods. The first is the PC regression proposed by Stock & Watson, (2002b); which we write as PC in our performance tables, PC-Squared (PC-Sq) and Squared-PC(Sq-PC) of Bai & Ng, (2008), kernel PCA (kPCA) [Kutateladze, (2022)], our linear counterpart, the 3PRF, and autoregressive model of lag order two [^8] We compared the performance across various autoregressive lags and found minimal differences. However, a lag of two generally showed superior performance in most cases.. Some of these methods require tuning of hyper-parameters to provide the best results, we do tune them as discussed in the subsection-5.1.2.
As discussed in section 2, different kernel functions correspond to different $\varphi(\cdot)$. We use the radial basis function (RBF) kernel.
$K_{\sigma}(\mathbf{x},\mathbf{x^{\prime}})=\exp\left(-\frac{\|\mathbf{x}-\mathbf{x^{\prime}}\|^{2}}{2\sigma^{2}}\right)$
where $\sigma$ is a hyperparameter determined via cross-validation. The use of this kernel is justified by its strong performance in macroeconomic forecasting, as documented by rbf_ref1, Exterkate et al., (2016), and Kutateladze, (2022).
@@ 5.1 Data and Hyper-Parameter Tuning
We utilize the quarterly macroeconomic dataset, FRED-QD. It spans the time period 1959-2023. This dataset encompasses a comprehensive array of more than 250 variables, including macroeconomic (such as GDP, Consumption, and Investment), financial, labor market, housing, and industrial and manufacturing variables. We present a tabulation comprising the mnemonic codes and details of the variables in the FRED-QD dataset alongside their counterparts in the Stock-Watson dataset in Supplementary appendix-B.4 for the series we forecast in this section.
@@ 5.1.1 Data Transformation
Estimation exercises involving nonstationary time series pose significant challenges. Nonstationarity is ubiquitous in economic and financial data. Nonstationary variables lack a defined population mean, and the sample standard deviation tends to diverge as the number of observations increases, see Onatski & Wang, (2021) for a more detailed discussion. Generally, researchers address this by manually examining each series to identify necessary transformations before computing principal components. Hamilton & Xi, (2024) offers an improved method for transforming the predictors to achieve stationarity. We use their method to make our data stationarity.
Scholars in the literature often employ sample periods devoid of structural breaks. Fan et al., (2023) notes that “There exist significant structural breaks for many variables around the year of the financial crisis in 2008 which makes our data non-stationary even after performing the suggested transformations”. Therefore, our study focuses on the stationary period spanning from 1965 to 2007999Another indirect advantage of the choice of this sample period is that it gives us the number of samples less than the number of predictors ($T<N$), hence a truly high-dimensional scenario to test our method in.. We conduct analyses on different combinations of the sample periods, including the entire available sample period from 1959 to 2023, and find no qualitative discrepancies in our findings.
In our main analysis, the sample period from 1964:Q4 to 2007:Q1 comprises $T=170$ observations (periods) and $N=176$ variables (predictors). While the data is initially available for around 250 series, those with missing values are removed; this leaves us with a total of 176 series. The model training and hyperparameter tuning are conducted within a rolling window framework, utilizing 70% of the total observations as the width of the rolling window. We observe qualitatively similar performance across varying window widths (50%, 60% of total data).
@@ 5.1.2 Hyper-parameter Tuning
Our methodology incorporates the kernel as a fundamental element of the estimation process. The kernel function includes a hyperparameter that necessitates optimization. Concurrently, a similar hyperparameter requires tuning in the context of a competitor method, namely kernel PCA. Thus, we employ an identical tuning procedure for both methodologies. We adopt a RBF kernel, which relies on a single hyperparameter, denoted as $\sigma$, for our specific applications. We partition the data into two folds and conduct cross-validation to determine the optimal tuning parameter. Further elaboration on the algorithm employed for this purpose is provided in the Supplementary appendix-B.5.
Furthermore, among our competitive methodologies, where factors are computed as PCs, we are required to specify the number of factors. To address this, we employ the eigenvalue ratio test method proposed by Ahn & Horenstein, (2013). This method computes the ratio of each eigenvalue to its predecessor and selects the number of principal components corresponding to the index where this ratio attains its maximum value. We employ a single factor throughout our analyses in both the 3PRF and kernel 3PRF models. This choice is often prudent within the 3PRF setting, as elucidated by Kelly & Pruitt, (2015), who highlight instances where a single factor can effectively represent a multi-factor system. When factors exhibit the same variances [^10] See appendix section A.7.2 in Kelly & Pruitt, (2015), a single proxy achieves optimal performance, and even when variances are not identical but closely aligned, one factor estimated through a single auto proxy typically explains a significant portion of the variation [^11] See simulations in appendix A.7.3 in Kelly & Pruitt, (2015), rendering residual variation minimal. While we assessed the performance of our estimator with varying numbers of factors, we consistently observed that a single factor predominates, thus, we report results based on this configuration.
@@ 5.2 Forecasting Using Theory Guided Proxies
The primary objective of this subsection is to establish the viability of theory-guided proxies in forecasting using our method. We also compare with the linear benchmark, i.e., 3PRF of Kelly & Pruitt, (2015). A more extensive performance evaluation will be presented in subsequent subsections, where we used the auto-proxy method discussed in table-3 to construct forecasts using Kernel 3PRF.
5.2.1 Forecasting GDP Using Investment and Consumption
We construct GDP forecasts using Consumption and investment as proxies and report the results in Table-4.
@@ 5.2.1 Forecasting GDP Using Investment and Consumption
Table 4: One-period Ahead Out-of-Sample $R^{2}$ for National Income
This exercise proves the efficacy of K3PRF over 3PRF while employing theory-guided proxies. Furthermore, our method outperforms the nearest competitive method Kelly & Pruitt, (2015).
5.2.2 Forecasting Inflation using Quantity Theory of Money
We reproduce the theory-guided proxy example discussed in Kelly & Pruitt, (2015). Δ(Price $\Delta(\text{Price level})$ i.e. inflation is our target variable for forecasting. The results for one-period ahead inflation forecasts are presented in Table-5. The quantity theory of money equation states that:
@@ 5.2.2 Forecasting Inflation using Quantity Theory of Money
Table 5: One-period Ahead Out-of-Sample $R^{2}$ for Inflation
The results indicate that the theory-guided proxies effectively capture inflation dynamics, yielding performance comparable to that of the closest competitor. It is important to emphasize again that this analysis focuses on one-step-ahead forecasts, which are not the primary strength of our methodology. The purpose of presenting these results is solely to demonstrate the workings of the procedure through the theory-guided proxies.
5.3 Comparative Forecasting Plots
To visually demonstrate the enhanced performance of kernel 3PRF compared to its linear counterpart, we provide comparative performance plots across four distinct types of economic series spanning various domains: macroeconomic series (Exports), price series (GDP Deflator), manufacturing series (Industrial Production), and financial series (S&P 500 Index) in figure-3 and 4. Plots of all other series on different forecast horizons are given in the Supplementary appendix-B.6.
@@ 5.3 Comparative Forecasting Plots
5.4 Forecasting Aggregate Macroeconomic Variables
@@
To present the results in an organized manner, we create two tables. In Table-6, we display the forecasting performance for three series: GDP, Consumption, and Investment, which we informally refer to as ‘Group-I’. Table-7 presents a comparative analysis of forecasting performance for ‘Group-II’ macro variables121212 Variables’ FRED-QD code and description can be found in Supplementary appendix-B.4: Exports, Imports, Fixed Investments, and Industrial Production (Final Index). As defined earlier in the text, the reported numbers in the tables represent out-of-sample $R^{2}$ values across various forecast horizons ranging from one period ahead to twelve periods ahead.
Results highlight a secular observation that among various unsupervised forecasting methodologies, PC, Squared-PC, PC-squared, and non-linear unsupervised approaches such as kernel PCA, none exhibit superior performance compared to our proposed method across any forecast horizon for the seven series under consideration. While the supervised linear forecasting model 3PRF demonstrates improved performance relative to the unsupervised techniques, it still falls short of outperforming our non-linear supervised approach. Notably, the autoregressive (AR) model emerges as the sole contender capable of surpassing our method in the shorter horizons, albeit only marginally and for a few series. However, our method significantly outperforms all the competitors across longer horizons. Therefore, our method emerges as a dependable and preferred forecasting framework across all forecast horizons in macroeconomic prediction tasks.
5.5 Forecasting Labor Market and Price Variables
This analysis aims to forecast key labor market and price variables. Within the labor market category, we focus on unemployment rates and total non-farm employment (Nonfarm Emp). We examine the GDP Deflator and the Consumer Price Index (CPI) for price variables. The GDP Deflator offers insights into overall inflation at the macroeconomic level, while the CPI captures inflation experienced by consumers at a more disaggregated level. The results of this analysis are summarized in Table 8.
Nonfarm Emp
@@ 5.5 Forecasting Labor Market and Price Variables
GDP Deflator
Method h=1 h=2 h=3 h=4 h=6 h=8 h=10 h=12
AR(2) 0.797 0.774 0.740 0.657 0.418 0.146 0.156 0.077
PC 0.444 0.276 0.056 -0.184 -0.408 -0.347 -0.221 -0.057
Sq-PC 0.299 0.145 -0.035 -0.168 -0.245 -0.230 -0.192 -0.108
PC-Sq 0.431 0.268 0.104 -0.039 -0.106 -0.038 -0.111 -0.182
kPCA -0.032 0.247 -0.021 0.008 0.003 0.004 0.029 -0.023
3PRF 0.584 0.496 0.426 0.243 0.174 0.279 0.300 0.155
k3PRF 0.667 0.632 0.563 0.476 0.479 0.413 0.197 0.512
CPI
Method h=1 h=2 h=3 h=4 h=6 h=8 h=10 h=12
AR(2) 0.704 0.706 0.620 0.565 0.397 0.211 0.062 -0.038
PC 0.660 0.535 0.364 0.154 -0.163 -0.252 -0.248 -0.173
Sq-PC 0.410 0.296 0.161 0.049 -0.055 -0.156 -0.200 -0.173
PC-Sq 0.649 0.512 0.353 0.186 -0.019 -0.087 -0.187 -0.228
kPCA 0.440 0.380 -0.050 0.189 -0.043 -0.024 0.042 -0.006
3PRF 0.641 0.566 0.487 0.352 0.192 0.241 0.255 0.141
k3PRF 0.676 0.612 0.541 0.463 0.469 0.434 0.349 0.477
Table 8: Out of Sample $R^{2}$ of Labor Market and Price Variables
Our relative forecast performance results are qualitatively similar to those of aggregate macroeconomic series forecasting.
5.6 Forecasting Housing and Financial Variables
We evaluate the relative performance of our method across several key indicators: Privately Owned Housing Starts (HStart), Privately Owned Housing Starts in the Western Census region (HStart-W), GS-1 (Treasury Bills), GS-10 (Treasury Notes), and the S&P 500 Index. The first two indicators pertain to the housing market, while the latter three belong to the financial market. These financial variables are listed in ascending order of volatility.